OpenStudy (anonymous):

How many electrons is 2415 Coulombs?

OpenStudy (anonymous):

oh...First you should understand that coulombs is related to Current*time: q=I*t So you ought to consider the time(s)^-^

OpenStudy (jamesj):

No, that's wrong. Coulomb is a measure of charge and there is no need to relate it to current. (Instead, we define current in terms of charge.) The charge on an electron is \[ e = 1.602 \times 10^{-19} \] Hence the number of electrons required for a charge of 2415 C is \[ \frac{2415}{e} = \frac{2415}{1.602 \times 10^{-19} = ... \]

OpenStudy (jamesj):

No, that's wrong. Coulomb is a measure of charge and there is no need to relate it to current. (Instead, we define current in terms of charge.) The charge on an electron is \[ e=1.602 \times 10^{−19} \] Hence the number of electrons required for a charge of 2415 C is \[ \frac{2415}{e} = \frac{2415}{1.602 \times 10^{-19}} = ... \]

OpenStudy (anonymous):

...that's right. sorry i mixed up